If x+y+z=1 and x^2+y^2+z^2=3 then find the value of xy+xz+yz. I really don't know where to begin. Any help that you could give me would be appreciated.
Answer by palanisamy(496)

Given, x+y+z=1 and x^2+y^2+z^2=3
Now, (x+y+z)^2 = x^2+y^2+z^2+2(xy+xz+yz)
1^2 = 3+2(xy+yz+zx)
1-3 = 2(xy+yz+zx)
-2 = 2(xy+yz+zx)
-2/2 = xy+yz+zx
xy+yz+zx = -1